A) A football is punted into the air. It has an initial vertical velocity of 15 m/s when it leaves the kicker’s foot (assume that air resistance is negligible).

What is the vertical velocity 1, 1.5, and 2 seconds after it leaves the kicker’s foot?

What is the vertical position 1, 1.5, and 2 seconds after it leaves the kicker’s foot?  

B) The great Black Jack Morris could pitch a 97 mph fastball. This is 142.2 ft/s. Distance from the mound to home plate is 60.5 ft. The ball is released 2.7 ft in front of the mound.

How much time does the batter have to react to this pitch?

C) Riggins runs the football 82 yds downfield before he is tackled. The straight-line distance between where he began the run and where he was tackled is 71 yd. The line makes an angle of 20° with the right side line.

How many yards did Riggins gain in this play?
Answer #1

A )

given

initial vertical velocity of 15 m/s

t = 1 sec

a = - g

y = u t + 0.5 a t2

y = ( 15 x 1 ) - 0.5 x 9.8 x 12

y = 10.1 m

t = 1.5 sec

a = - g

y = u t + 0.5 a t2

y = ( 15 x 1.5 ) - 0.5 x 9.8 x 1.52

y = 11.475 m

t = 2 sec

a = - g

y = u t + 0.5 a t2

y = ( 15 x 2 ) - 0.5 x 9.8 x 22

y = 10.4 m

B )

using time = distance / speed

distance = 60.5 - 2.7 = 57.8 ft

time = 57.8 / 142.2

t = 0.406 sec

C )

d = 71 m

\theta = 20o

dx = 71 cos20

= 66.72 yards

dy = 71 sin20

= 24.28 yards

number of yards gained by Riggins in this play is = 66.72 yards